\begin{thebibliography}{99}







\bibitem{Adams-75}
R. A. Adams, \textit{Sobolev spaces}, 
Academic Press, Boston, MA, USA, 1975.

\bibitem{Benguria-79} R. Benguria, {\it The von Weizs\"acker and exchange corrections in the Thomas--Fermi theory}, PhD. Thesis, Princeton University, 1979.

\bibitem{BBL-81} R. Benguria, H. Brezis, and E. H. Lieb, ``The Thomas--Fermi--von Weizs\"acker theory of atoms and molecules'', {\it Commun. Math. Phys.} {\bf 79}, 167--180 (1981).



\bibitem{Besicovitch-45} A. S. Besicovitch, ``A general form of the covering principle and relative differentiation of additive functions", {\em Proc. Cambridge Philos. Soc.} {\bf 41}, 103--110 (1945).%; {\bf 42}, 205--235 (1946). 




%
%\bibitem{AizLie-78} 
%M. Aizenman and  E. H. Lieb, 
%``On semi-classical bounds for eigenvalues of Schr\"odinger operators'', 
%\textit{Phys.\ Lett.\ A\/} \textbf{66}, 427--429 (1978).
%
%\bibitem{BelGhiMerMorSch-18}
%J. Bellazzini, M. Ghimenti, C. Mercuri, V. Moroz,
%and J. Van Schaftingen
%``Sharp Gagliardo--Nirenberg inequalities in fractional Coulomb--Sobolev
%spaces'', 
%\textit{Trans.\ Amer.\ Math.\ Soc.\/} \textbf{370}, 8285--8310 (2018).
%
%\bibitem{BelOzaVis-11}
%J. Bellazzini, T. Ozawa, and N. Visciglia,
%``Ground states for semi-relativistic Schr{\"o}dinger--Poisson--Slater
%energies'',
%e-print arXiv:1103.2649 (2011).
%
%\bibitem{BelFraVis-14}
%J. Bellazzini, R. L. Frank, and N. Visciglia,
%``Maximizers for Gagliardo--Nirenberg inequalities and related non-local
%problems'',
%\textit{Math.\ Ann.\/} \textbf{360}, 653--673 (2014).
%
%\bibitem{BenBleLos-12}
%R. D. Benguria, G. A. Bley, and M. Loss,
%``A new estimate on the indirect Coulomb energy'',
%\textit{Int.\ J. Quantum Chem.\/} \textbf{112}, 1579--1584 (2012).
%
%\bibitem{BlaStu-96}
%P. Blanchard and J. Stubbe,
%``Bound states for Schr\"odinger Hamiltonians: Phase Space Methods and
%Applications'',
%\textit{Rev.\ Math.\ Phys.\/} \textbf{35}, 504--547 (1996).  
%
%\bibitem{BleFou-19}
%G. A. Bley and S. Fournais,
%``Hardy--Lieb--Thirring inequalities for fractional Pauli operators'',
%\textit{Commun.\ Math.\ Phys.\/} \textbf{365}, 651--683 (2019). 
%
%\bibitem{CheXu-97} 
%J.-Y. Chemin and C.-J. Xu, 
%``Inclusions de Sobolev en calcul de Weyl-H\"ormander et champs de vecteurs
%sous-elliptiques'', 
%\textit{Ann.\ Sci.\ \'Ecole Norm.\ Sup.\/} \textbf{30}, 719--751 (1997).
%
\bibitem{Cwikel-77}
M. Cwikel,
``Weak type estimates for singular values and the number of bound states of
Schr\"odinger operators'',
\textit{Annals of Mathematics} \textbf{106}, 93--100 (1977). 
%
%\bibitem{Daubechies-83}
%I. Daubechies, ``An uncertainty principle for fermions with generalized kinetic
%energy'', \textit{Commun.\ Math.\ Phys.\/} \textbf{90}, 511--520 (1983).
%
\bibitem{DolLapLos-08}
J. Dolbeault, A. Laptev, and M. Loss,
``Lieb--Thirring inequalities with improved constants'',
\textit{J. Eur.\ Math.\ Soc.\/} \textbf{10},  1121--1126 (2008).
%
\bibitem{DysLen-67}
F. J. Dyson and A. Lenard,
``Stability of matter. I'',
\textit{J. Math.\ Phys.\/} \textbf{8},  423--434 (1967).

\bibitem{DysLen-67b}
F. J. Dyson and A. Lenard, ``Stability of matter. II'',
\textit{J. Math.\ Phys.\/} \textbf{9},  698--711 (1968).

%\bibitem{EkhFra-06}
%T. Ekholm and R. L. Frank,
%``On Lieb--Thirring inequalities for Schr\"odinger operators with virtual
%level'',
%\textit{Commun.\ Math.\ Phys.\/}, \textbf{264}, 725--740 (2006).
%
\bibitem{EdeFoi-91}
  A. Eden and C. Foias,
  ``A simple proof of the generalized Lieb--Thirring
  inequalities in one-space dimension'',
  \textit{J. Math.\ Anal.\ Appl.\/} \textbf{162}, 250--254 (1991).
%
%\bibitem{Frank-09}
%  R. L. Frank,
%  ``A simple proof of Hardy--Lieb--Thirring inequalities'',
%  \textit{Commun.\ Math.\ Phys.\/}, \textit{290}, 789--800 (2009).
%
%\bibitem{Frank-14}
%  R. L. Frank,
%  ``Cwikel's theorem and the CLR inequality'',
%  \textit{J. Spectral Theory\/} \textbf{4}, 1--21 (2014).
%
%\bibitem{Frank-18}
%  R. L. Frank,
%  ``Eigenvalue bounds for the fractional Laplacian: a review'',
%  In: \textit{Recent development in nonlocal theory\/}, 
%  De Gruyter, Berlin, Germany, 2018, pp.~210--235.
%
\bibitem{Frank-20}
  R. L. Frank,
  ``The Lieb--Thirring inequalities: Recent results and open problems'', In: {\em Nine mathematical challenges: an elucidation}, A. Kechris, et al. (eds.), 45 - 86, Proceedings of Symposia in Pure Mathematics 104, Amer. Math. Soc., Providence, RI, 2021. arXiv:2007.09326

%\bibitem{FraGonLew-20}
%  R. L. Frank, D. Gontier, and M. Lewin,
%  ``The nonlinear Schr\"odinger equation for orthonormal functions:
%  II. Application to Lieb--Thirring inequalities'',
%  \textit{Commun.\ Math.\ Phys.\/} (to appear), eprint arXiv:2002.04964 (2020).
%
%\bibitem{FraHunJexNam-18}
%  R. L. Frank, D.  Hundertmark, M. Jex and P. T. Nam,
%  ``The Lieb--Thirring inequality revisited'',
%  \textit{J. Eur.\ Math.\ Soc.\/} \textbf{23}, 2583--2600 (2021).
%
%\bibitem{FraLieSei-07}
%  R. L. Frank, E. H. Lieb, and R. Seiringer,
%  ``Hardy--Lieb--Thirring inequalities for fractional Schr\"odinger
%  operators'',
%  \textit{J. Amer.\ Math.\ Soc.\/} \textbf{21}, 925--950 (2007).
%
%\bibitem{FraLewLieSei-11}
%  R. L. Frank, M. Lewin, E. H. Lieb, and R. Seiringer,
%  ``Energy Cost to Make a Hole in the Fermi Sea'',
%  \textit{Phys.\ Rev.\ Lett.\/} \textbf{106}, 150402 (2011).
%
%\bibitem{FraLewLieSei-13}
%  R. L. Frank, M. Lewin, E. H. Lieb, and R. Seiringer,
%  ``A positive density analogue of the Lieb--Thirring inequality'',
%  \textit{Duke Math.\ J.} \textbf{162}, 435-495 (2013).
%
%\bibitem{FraSei-12}
%  R. L. Frank and R. Seiringer,
%  ``Lieb--Thirring inequality for a model of particles with point
%  interactions'',
%  \textit{J. Math.\ Phys.\/} \textbf{53},  095201 (2012).
%
%\bibitem{GotNam-18}
%  N. Gottschling and P. T. Nam,
%  ``Convergence of Levy--Lieb to Thomas--Fermi density functional'', 
%   \textit{Cal.\ Var.\ PDE\/} \textbf{57}, 105--117 (2018).
%
%\bibitem{Herbst-77}
%  I. W. Herbst,
%  ``Spectral theory of the operator $(p^2+m^2)^{1/2}-Ze^2/r$'',
%  \textit{Commun.\ Math.\ Phys.\/} \textbf{53}, 285--294 (1977).
%
\bibitem{HO-77} 
  M. Hoffmann-Ostenhof and T. Hoffmann-Ostenhof,
  ``\,``Schr\"odinger inequalities'' and asymptotic behavior of the electron density
  of atoms and molecules'',
  \textit{Phys.\ Rev.\ A} \textbf{16},  1782--1785 (1977).

%\bibitem{HunLapWei-00}
%  D. Hundertmark, A. Laptev, and T. Weidl,
%  ``New bounds on the Lieb--Thirring constant'',
%  \textit{Invent.\ Math.\/} \textbf{140}, 693--704 (2000).
%
%\bibitem{HunLieTho-98}
%  D. Hundertmark, E. H. Lieb, and L. E. Thomas,
%  ``A sharp bound for an eigenvalue moment of the one-dimensional
%  Schr\"odinger operator'',
%  \textit{Adv.\ Theor.\ Math.\ Phys.\/} \textbf{2}, 719--731 (1998). 
%
%\bibitem{HunKunRieVug-18}
%  D. Hundertmark, P. Kunstmann, T. Ried, and S. Vugalter,
%  ``Cwikel's bound reloaded'', e-print arXiv:1809.05069 (2018).
%
%\bibitem{IlyLapZel-20}
%  A. Ilyin, A. Laptev, and S. Zelik,
%  ``Lieb--Thirring constant on the sphere and on the torus'',
%  \textit{J. Funct.\ Anal.\/} \textbf{279}, 108784 (2020). 
%
\bibitem{KogNam-20}
  K. K\"ogler and P. T. Nam,
  ``The Lieb--Thirring inequality for interacting systems in strong-coupling
  limit'',
  \textit{Arch.\ Rational Mech.\ Anal.} \textbf{240}, 1169--1202 (2021).
%
%\bibitem{LapWei-00}
%  A. Laptev and T. Weidl,
%  ``Sharp Lieb--Thirring inequalities in high dimensions'',
%  \textit{Acta Math.\/} \textbf{184}, 87--111 (2000).
%
%\bibitem{LarLun-18}
%  S. Larson and D. Lundholm,
%  ``Exclusion Bounds for Extended Anyons'',
%  \textit{Arch.\ Rational Mech.\ Anal.\/} \textbf{227}, 309--365 (2018). 
%
%\bibitem{LarLunNam-19}
%  S. Larson, D. Lundholm, and P. T. Nam,
%  ``Lieb--Thirring inequalities for wave functions vanishing on the diagonal
%  set'',
%  \textit{Annales Henri Lebesgue\/} \textbf{4}, 251--282 (2021). 
%
%\bibitem{Levy-79}
%  M. Levy,
%  ``Universal variational functionals of electron densities, first-order density
%  matrices, and natural spin-orbitals and solution of the v-representability
%  problem'',
%  \textit{Proc.\ Natl.\ Acad.\ Sci.\ U.\ S.\ A.\/} \textbf{76}, 6062--6065 (1979).
%
%\bibitem{LewLie-15}
%  M. Lewin and E. H. Lieb,
%  ``Improved Lieb--Oxford exchange-correlation inequality with gradient
%  correction'',
%  \textit{Phys.\ Rev.\ A\/} \textbf{91}, 022507 (2015). 
%
%\bibitem{LewLieSei-20}
%  M. Lewin,  E. H. Lieb, and R. Seiringer,
%  ``The local density approximation in density functional theory'',
%  \textit{Pure Appl.\ Analysis\/} \textbf{2}, 35--73 (2020). 
%
%\bibitem{LewLieSei-19}
%  M. Lewin,  E. H. Lieb, and R. Seiringer,
%  ``Universal Functionals in Density Functional Theory'',
%  In: \textit{Density Functional Theory\/}, edited by E. Canc\`es, G.
%  Friesecke, and L. Lin, e-print arXiv:1912.10424 (2019). 
%
\bibitem{Lieb-76}
  E. H. Lieb,
  ``Bounds on the eigenvalues of the Laplace and Schr\"odinger operators'',
  \textit{Bull.\ Am.\ Math.\  Soc.\/} \textbf{82}, 751--54 (1976). 
%
%\bibitem{Lieb-83b}
%  E. H. Lieb,
%  ``An $L^p$ bound for the Riesz and Bessel potentials of orthonormal
%  functions'',
%  \textit{J. Funct.\ Anal.\/} \textbf{51}, 159--165 (1983).
%
%\bibitem{Lieb-83}
%  E. H. Lieb,
%  ``Density functionals for Coulomb systems'',
%  \textit{Int.\ J. Quantum Chem.\/} \textbf{24}, 243--277 (1983).
%
%\bibitem{Lieb-84}
%  E. H. Lieb,
%  ``On characteristic exponents in turbulence'',
%  \textit{Commun.\ Math.\ Phys.\/} \textbf{82}, 473--480 (1984).
%
%\bibitem{Lieb-00}
%  E. H. Lieb,
%  ``Lieb--Thirring Inequalities'',
%  in \textit{Encyclopedia of Mathematics, Supplement vol.~II\/},
%  Kluwer, Dordrecht, The Netherlands, 2000; pp.~311--313.  
%
\bibitem{LieLos-01}
  E. H. Lieb and M. Loss,
  \textit{Analysis\/}, second edition,
    Graduate Studies in Mathematics,
    American Mathematical Society, Providence, RI, USA, 2001.
%
%\bibitem{LieOxf-80}
%  E. H. Lieb and S. Oxford,
%  ``Improved lower bound on the indirect Coulomb energy'',
%  \textit{Int.\ J. Quantum Chem.\/} \textbf{19}, 427--439 (1981).
%
\bibitem{LieSei-10}
  E. H. Lieb and R. Seiringer,
  \textit{The Stability of Matter in Quantum Mechanics\/},
  Cambridge University Press, Cambridge, UK, 2010. 

\bibitem{LieThi-75}
  E. H. Lieb and W. E. Thirring,
  ``Bound on kinetic energy of fermions which proves stability of matter'',
  \textit{Phys.\ Rev.\ Lett.\/} \textbf{35}, 687--689 (1975).

\bibitem{LieThi-76}
  E. H. Lieb and W. E. Thirring,
  ``Inequalities for the moments of the eigenvalues of the Schr\"odinger
  Hamiltonian and their relation to Sobolev inequalities'',
  in \textit{Studies in Mathematical Physics},
  Princeton University Press, %Princeton, NJ, USA, 
  1976;  pp.~269--303.


%\bibitem{Lundholm-18}
%  D. Lundholm,
%  ``Methods of modern mathematical physics: Uncertainty and exclusion
%  principles in quantum mechanics''  (lecture notes), e-print arXiv:1805.03063
%  (2018); the latest revision is available at
%  \url{http://www.mathematik.uni-muenchen.de/~lundholm/methmmp.pdf}. 
%
\bibitem{LunNamPor-16}
  D. Lundholm,  P. T. Nam, and F. Portmann,
  ``Fractional Hardy--Lieb--Thirring and related inequalities for interacting
  systems'',
  \textit{Arch.\ Rational Mech.\ Anal.\/} \textbf{219}, 1343--1382 (2016).

%\bibitem{LunPorSol-15}
%  D. Lundholm, F. Portmann, and J. P. Solovej,
%  ``Lieb--Thirring bounds for interacting Bose gases'',
%  \textit{Commun.\ Math.\ Phys.\/} \textbf{335}, 1019--1056 (2015).
%
%\bibitem{LunQva-20}
%  D. Lundholm and V. Qvarfordt,
%  ``Exchange and exclusion in the non-abelian anyon gas'',
%  e-print arXiv:2009.12709 (2020).
%
%\bibitem{LunSol-13a} 
%  D. Lundholm and J. P. Solovej,
%  ``Local exclusion principle for identical particles obeying intermediate and
%  fractional statistics'', 
%  \textit{Phys.\ Rev.\ A \/} \textbf{88}, 062106 (2013). 

\bibitem{LunSol-13}
  D. Lundholm and J. P. Solovej,
  ``Hardy and Lieb--Thirring inequalities for anyons'',
  \textit{Commun.\ Math.\ Phys.\/} \textbf{322}, 883--908 (2013).

\bibitem{LunSol-14}
  D. Lundholm and J. P. Solovej,
  ``Local exclusion and Lieb--Thirring inequalities for intermediate and
  fractional statistics'',
  \textit{Ann.\ Henri Poincar\'e\/} \textbf{15}, 1061--1107 (2014).

%\bibitem{LunSei-18}
%  D. Lundholm and R. Seiringer. 
%  ``Fermionic behavior of ideal anyons'',
%  \textit{Lett.\ Math.\ Phys.\/} \textbf{108}, 2523--2541 (2018).
%
%\enlargethispage{-0.2\baselineskip}%%++
%  
%\bibitem{MarYou-58}
%  N. H. March and W. H. Young.
%  ``Variational Methods based on the Density Matrix'',
%  \textit{Proc.\ Phys.\ Soc.\/} \textbf{72}, 182--192 (1958). 
%
\bibitem{MalNgu-21} A. Mallick and H.-M. Nguyen, ``Gagliardo-Nirenberg and Caffarelli-Kohn-Nirenberg interpolation inequalities associated with Sobolev-Coulomb spaces", \textit{J. Funct. Anal.} {\bf 283}, 109662 (2022).

\bibitem{Morse-47} A.P. Morse, ``Perfect blankets'', {\it Trans. Amer. Math. Soc.} {\bf  61}, 418--442  (1947). 




%\bibitem{Nam-18}
%  P. T. Nam,
%  ``Lieb--Thirring inequality with semiclassical constant and gradient error
%  term'',
%  \textit{J. Funct.\ Anal.\/} \textbf{274}, 1739--1746 (2018). 

\bibitem{Nam-19}
  P. T. Nam, ``Direct methods to Lieb-Thirring kinetic inequalities", In: {\em Proceedings of the Workshop on Density Functionals for Many-Particle Systems}, September 2019, Singapore. arXiv:2012.12045

%\bibitem{NetWei-96}
%  Y. Netrusov and T. Weidl,
%  ``On Lieb--Thirring inequalities for higher order operators with critical and
%  subcritical powers'',
%  \textit{Commun.\ Math.\ Phys.\/} \textbf{182}, 355--370 (1996).
%
\bibitem{Rozenblum-76}
  G. V. Rozenblum,
  ``Distribution of the discrete spectrum of singular differential
  operators'',
  \textit{Izvestiya Vysshikh Uchebnykh Zavedenii Matematika\/} \textbf{1},
  75--86 (1976). 
  
\bibitem{Rumin-10}
  M. Rumin,
  ``Spectral density and Sobolev inequalities for pure and mixed states'',
  \textit{Geom.\ Funct.\ Anal.\/} \textbf{20}, 817--844 (2010).

\bibitem{Rumin-11}
  M. Rumin,
  ``Balanced distribution-energy inequalities and related entropy bounds'',
  \textit{Duke Math.\ J.\/} \textbf{160}, 567--597 (2011).

\bibitem{Sabin-16}
  J. Sabin,
  ``Littlewood--Paley decomposition of operator densities and application to a
  new proof of the Lieb--Thirring inequality'',
  \textit{Math.\ Phys.\ Anal.\ Geom.\/} \textbf{19}, 11 (2016).

%\bibitem{Simon-05}
%  B. Simon,
%  \textit{Trace Ideals and Their Applications\/},
%  Mathematical Surveys and Monographs Vol.~120, second edition,
%  American Mathematical Society, 2005.
%
%\bibitem{SolSorSpi-10}
%  J. P. Solovej, T. \O. S\o rensen, and W. L. Spitzer,
%  ``Relativistic Scott correction for atoms and molecules'',
%  \textit{Comm.\ Pure Appl.\ Math.\/} \textbf{63}, 39--118 (2010). 
%
%\bibitem{Weidl-96}
%  T. Weidl,
%  ``On the Lieb--Thirring constants $L_\gamma$ for $\gamma \ge 1/2$'',
%  \textit{Commun.\ Math.\ Phys.\/} \textbf{178}, 135--146 (1996).  
%
%\bibitem{Weyl-11}
%  H. Weyl,
%  ``\"Uber die asymptotische Verteilung der Eigenwerte'',
%  \textit{Nachrichten der K\"oniglichen Gesellschaft der Wissenschaften zu
%    G\"ottingen\/} 110--117 (1911). 
%
%\bibitem{Weyl-12}
%  H. Weyl,
%  ``Das asymptotische Verteilungsgesetz linearen partiellen
%  Differentialgleichungen'',
%  \textit{Math.\ Ann.\/} \textbf{71}, 441--479 (1912). 

\end{thebibliography}

\documentclass[11pt,a4paper,reqno]{amsart}%
\usepackage{amsthm,amsmath,amsfonts,amssymb,amsxtra,appendix,bookmark,dsfont,bm,mathrsfs,color}
\usepackage{bbm} 

\usepackage[a4paper,left=40mm,top=25mm,bottom=25mm,right=40mm]{geometry}

\newtheorem{theorem}{Theorem}

\newtheorem{lemma}[theorem]{Lemma}

\theoremstyle{definition}

\theoremstyle{remark}

\DeclareMathOperator{\Tr}{Tr}

\def\bq{\begin{eqnarray}}
\def\eq{\end{eqnarray}}
\def\bqq{\begin{align*}}
\def\eqq{\end{align*}}

\def\eps{\varepsilon}

\newcommand\1{{\ensuremath {\mathds 1} }}

\def\R {\mathbb{R}}

\def\R {\mathbb{R}}

\def\d{{\rm d}}

\title[] {A proof of the Lieb--Thirring inequality via the Besicovitch covering lemma}

\author[P. T. Nam]{Phan Th\`anh Nam} 
\address{LMU Munich, Department of Mathematics, Theresienstrasse 39, D-80333 Munich, Germany} 
\email{nam@math.lmu.de}

\begin{document}

\maketitle

\begin{center}
{\em Dedicated to Professor Duong Minh Duc on the occasion of his 70th birthday}  
\end{center}

\begin{abstract} We give a proof of the Lieb--Thirring inequality on the kinetic energy of orthonormal functions by using a microlocal technique, in which the  uncertainty and exclusion principles are combined through the use of the Besicovitch covering lemma. 

%
\end{abstract}

\date{\today}


\section{Introduction}

The celebrated Lieb-Thirring inequality \cite{LieThi-75,LieThi-76} states that for every trace class operator $0\le \gamma \le 1$ on $L^2(\R^d)$ with density $\rho_\gamma(x)=\gamma(x,x)$, the kinetic energy lower bound 
\bq \label{eq:LT}
\Tr(-\Delta \gamma) \ge K_d \int_{\R^d} \rho_\gamma(x)^{1+\frac{2}{d}} \d x
\eq
holds with a  constant $K_d>0$ depending only on the dimension $d\ge 1$ (in particular, $K_d$ is independent of $\gamma$). Here assuming the spectral decomposition 
\bq \label{eq:gamma-un} \gamma= \sum_{n=1}^\infty \lambda_n |u_n\rangle \langle u_n|
\eq
with $\{u_n\}_{n=1}^\infty$ being orthonormal in $L^2(\R^d)$ and  $0\le \lambda_n\le 1$, we denote  
\bq \label{eq:gamma-rho}
\Tr(-\Delta \gamma)  = \sum_{n=1}^\infty \lambda_n \int_{\R^d} |\nabla u_n|^2, \quad \rho_\gamma(x)=\sum_{n=1}^\infty \lambda_n |u_n(x)|^2. 
\eq

Strictly speaking, the trace class condition $\Tr \gamma=\sum_{n=1}^\infty \lambda_n <\infty$ is not essential. The bound \eqref{eq:LT} still holds for every operator $0\le \gamma\le 1$ provided that $\gamma$ admits a spectral decomposition as in \eqref{eq:gamma-un} such that we can define the kinetic energy $\Tr(-\Delta \gamma)$ and the density $\rho_\gamma$ via \eqref{eq:gamma-rho} (this extension can be obtained easily from a density argument using the Lebesgue monotone convergence theorem). In connection to quantum physics, $\gamma$ usually stands for the one-body density matrix of a quantum state, and the trace class condition means that the number of particles is finite. A  key feature of \eqref{eq:LT} is that the bound is independent of the number of particles, thus allowing us to treat large systems (including even infinite systems via a standard limiting argument).

By some sort of the uncertainty principle (e.g. Sobolev's inequality), it is not difficult to see that
\bq \label{eq:S}
\left( \int_{\R^d} |\nabla u|^2 \right) \left( \int_{\R^d} |u|^2 \right)^{\frac 2 d} \ge K_{\rm So} \int_{\R^d} |u(x)|^{2(1+\frac{2}{d})} \d x, \quad \forall u\in H^1(\R^d).
\eq
Here the optimal constant $K_{\rm So}>0$ depends only on the dimension $d\ge 1$. Consequently, if $\{u_n\}_{n=1}^N$ is a family of {\em normalized} functions in $L^2(\R^d)$, then using \eqref{eq:S} we have
\bq \label{eq:S1}
\sum_{n=1}^N \int_{\R^d} |\nabla u_n|^2 \ge   K_{\rm So} \int_{\R^d} \sum_{n=1}^N |u_n|^{2(1+\frac{2}{d})}   \ge \frac{K_{\rm So} }{N^{2/d}}  \int_{\R^d} \left( \sum_{n=1}^N |u_n|^{2}  \right)^{1+\frac 2 d}. 
\eq 
The optimality of the constant $K_{\rm So} N^{-2/d}$ in \eqref{eq:S1} can be seen easily by considering the special case $u_n=u$ for all $n$. On the other hand, if $\{u_n\}_{n=1}^N$ are {\em orthonormal}, then the Lieb-Thirring inequality \eqref{eq:LT} with $\gamma=\sum_{n=1}^N |u_n\rangle \langle u_n|$ implies that  \eqref{eq:S1} holds with $K_{\rm So} N^{-2/d}$ replaced by a constant $K_d$ depending only on the dimension $d\ge 1$, which is very useful when $N\to \infty$. The orthogonality here essentially corresponds to the condition $0\le \gamma\le 1$ in \eqref{eq:LT}.

As an application, the Lieb--Thirring inequality \eqref{eq:LT} gives a semiclassical estimate for the kinetic energy of fermionic systems, which is a crucial ingredient of the proof of the stability of matter in \cite{LieThi-75}. In this aspect, the constraint $0\le \gamma \le 1$ corresponds to Pauli's exclusion principle of fermions (see e.g.  \cite[Theorem 3.2]{LieSei-10}). Therefore,  \eqref{eq:LT} can be interpreted as an elegant combination of the uncertainty and exclusion principles in quantum mechanics.

Historically, in the original proof of Lieb and Thirring in  \cite{LieThi-75,LieThi-76}, the kinetic bound \eqref{eq:LT} was derived from its dual version 
\bq \label{eq:LT-V}
\Tr[(-\Delta + V(x))_-]  \ge -L_d \int_{\R^d} |V_-(x)|^{1+\frac d 2} \d x
\eq
where $t_-=\min(t,0)$ and 
$$
K_d\left( 1 + \frac 2 d \right) = \left[ L_d \left( 1+ \frac d 2\right) \right]^{-\frac 2 d}.
$$
In this approach, the sum of all negative eigenvalues of the Schr\"odinger operator $-\Delta + V(x)$ is estimated via the number of eigenvalues $\ge 1$ of the
Birman--Schwinger operator  $
\sqrt{|V_-(x)|} (-\Delta +E)^{-1} \sqrt{|V_-(x)|}
$ with $E>0$. This duality approach is powerful since it relates the nonlinear inequality \eqref{eq:LT} to a more linear problem, namely a spectral property of a linear operator.  The spectral estimate \eqref{eq:LT-V} is interesting in its own right and has been generalized in various directions; see \cite{Frank-20} for a recent review with many open questions. 

In this article, we will give a direct proof of the kinetic bound \eqref{eq:LT}, without relying on the duality argument. The first proof in this direction goes back to the work of Eden and Foias in 1991 \cite{EdeFoi-91} for one dimension, which was later extended by Dolbeault, Laptev, and Loss \cite{DolLapLos-08} for higher dimensions. More recently, two other direct proofs were given by Rumin \cite{Rumin-10,Rumin-11} and by Lundholm and Solovej \cite{LunSol-13,LunSol-14}; we refer to \cite{Nam-19} for a review on these two approaches. See also \cite{Sabin-16} for another direct proof. 

We will discuss a simplification of the Lundholm--Solovej approach. The key observation of this approach, which actually goes back to the first proof of the stability of matter by Dyson and Lenard \cite{DysLen-67,DysLen-67b},  is that a local version of the Lieb--Thirring inequality on cubes can be obtained easily by local versions of the uncertainty and exclusion principles. As realized in \cite{LunSol-13,LunSol-14}, it is possible to obtain the global bound  \eqref{eq:LT} by combining such local estimates by choosing an appropriate family of disjoint cubes. A simplification of the Lundholm--Solovej approach was already given by Lundholm--Nam--Portmann \cite{LunNamPor-16} where the local uncertainty and exclusion principles are combined by using a new covering lemma which is inspired by a ``stopping time argument" in harmonic analysis. Our aim here is to provide with a further simplification where the local uncertainty and exclusion principles are combined naturally via the classical {\em Besicovitch covering lemma}, thus making the proof very elementary. We hope that our proof will be not only interesting from the pedagogical point of view, but also helpful for finding new generalizations of \eqref{eq:LT}   in the future. In particular, we expect that the method represented here can be strengthened by coupling with the ideas of using dyadic sets in recent works \cite{KogNam-20,MalNgu-21}.

Note that the Besicovitch covering lemma was already used in Rozenblum's proof of the Cwikel--Lieb--Rozenblum inequality \cite{Cwikel-77,Lieb-76,Rozenblum-76} on the number of negative eigenvalues of Schr\"odinger operators for $d\ge 3$:
\bq \label{eq:CLR-V}
\Tr \Big[ \1_{(-\infty,0)} ( -\Delta + V(x) ) \Big] \le C_d \int_{\R^d} |V_-(x)|^{\frac d 2} \d x. 
\eq
In this approach, using Sobolev's inequality, the linear operator $-\Delta + V(x)$ can be localized in bounded domains such that it has at most one bound state in each domain, and then the number of the domains is controlled by the Besicovitch covering lemma. Heuristically, since \eqref{eq:CLR-V} is stronger than \eqref{eq:LT-V} (when $d\ge 3$), it is not surprising that the Besicovitch covering lemma is also helpful for the proof of the Lieb--Thirring inequality. However, since \eqref{eq:LT}  seems less linear than \eqref{eq:LT-V} and \eqref{eq:CLR-V}, we think it is still interesting to see that a variant of this microlocal technique works directly for \eqref{eq:LT}. 

%

\section{Proof}

Our proof of \eqref{eq:LT} includes two steps. In the first step, we derive a local version of \eqref{eq:LT} in any ball $B\subset \R^d$ such that  $\int_B \rho_\gamma =2$ (we can replace $2$ by any fixed number bigger than 1). This step is done by combining local versions of the uncertainty and exclusion principles, exactly in the spirit of the proofs in \cite{LunSol-13,LunSol-14,LunNamPor-16} (the only difference is that we are interested in balls rather than cubes). In the second step, we cover $\R^d$ by a suitable collection of balls via the Besicovitch covering lemma, and conclude   \eqref{eq:LT}.

Now we come to the details. Let us denote, with the spectral decomposition $\gamma= \sum_n \lambda_n |u_n\rangle \langle u_n|$ and any ball $B\subset \R^d$, that  
\begin{align} \label{eq:Tr-B-local}
\Tr(-\Delta_B \gamma)  = \sum_n \lambda_n  \int_{B}  |\nabla u_n|^2, \quad \rho_\gamma(x)=\sum_n \lambda_n  |u_n(x)|^2. 
\end{align}
We will always denote by $C_d>0$ a general (large) constant depending only on the dimension $d\ge 1$. The value of $C_d$ may change line by line. 

\bigskip
\noindent
{\bf Step 1:} First, we prove a local version of \eqref{eq:LT}. 

\begin{lemma}[Lieb--Thirring inequality for balls] \label{lem:LT-ball}For any ball $B\subset \R^d$ and any trace class operator $0\le \gamma \le 1$ on $L^2(B)$ satisfying $\int_B \rho_\gamma = 2$, we have
\bq \label{eq:LT-ball}
\Tr(-\Delta_B \gamma) \ge \frac{1}{C_d} \int_{B} \rho_\gamma(x)^{1+\frac{2}{d}} \d x.
\eq
The constraint $\int_B \rho_\gamma = 2$ can be replaced by $a\le \int_B \rho_\gamma \le b$ for any constants $1<a<b<\infty$ (then $C_d$ depends on $d,a,b$). 
\end{lemma}

Lemma \ref{lem:LT-ball} is a consequence of the following well-known results. 

\begin{lemma}[Hoffmann-Ostenhof inequality] \label{lem:HO} For any measurable subset $B\subset \R^d$ and any trace class operator $0\le \gamma \le 1$ on $L^2(B)$ we have
$$
\Tr(-\Delta_B \gamma)  \ge \int_B |\nabla \sqrt{\rho_\gamma}|^2. 
$$
\end{lemma}
\begin{proof} This result was first proved in \cite{HO-77}. For the reader's convenience, here we recall another derivation using the diamagnetic inequality $|\nabla |u| | | \le |\nabla u|$ (see \cite[Theorem 7.21]{LieLos-01}). Indeed, for two complex functions $\{v_n\}_{n=1}^2$, using the diamagnetic inequality for $v_1, v_2$ and $|v_1|+ {\bf i} |v_2|$ (with ${\bf i}^2=-1$) we have 
\begin{align} \label{eq:u1u2}
|\nabla v_1|^2 + |\nabla v_2|^2 &\ge |\nabla |v_1||^2 + |\nabla |v_2||^2 = |\nabla (|v_1| + {\bf i} |v_2|)|^2 \ge \left |\nabla \sqrt{ |v_1|^2 + |v_2|^2}\right|^2.  
\end{align}
Note that \eqref{eq:u1u2} is called the convexity of the gradient which was first published in \cite[Theorem 4.3]{Benguria-79} and \cite[Lemma 4]{BBL-81} (see also \cite[Theorem 7.8]{LieLos-01}). By induction, we find that for any functions $\{v_n\}_{n=1}^\infty$,
\begin{align} \label{eq:u1u2un}
\sum_{n=1}^\infty |\nabla v_n|^2 \ge \left|\nabla \sqrt{ \sum_{n=1}^\infty |v_n|^2}\right|^2.  
\end{align}
We obtain the desired estimate by integrating the pointwise inequality \eqref{eq:u1u2un} with $v_n=\sqrt{\lambda_n} u_n$ and using the definition \eqref{eq:Tr-B-local}. 
\end{proof}

\begin{lemma}[Local uncertainty principle] \label{lem:LU} For any ball $B\subset \R^d$ and any function $g\in H^1(B)$ we have 
$$
\int_B |\nabla g|^2 \ge \frac{1}{C_d} \frac{\int_B |g|^{2(1+ \frac 2 d)}}{(\int_B |g|^2)^{\frac 2 d}} - \frac{C_d}{ |B|^{\frac 2 d}} \int_B |g|^2 .
$$
\end{lemma}

\begin{proof} By scaling we may assume that $B$ is the unit ball. Then by a standard result (see \cite[Theorem 7.41]{Adams-75}), we can find 
$G\in H^1(\R^d)$, an extension of $g$, such that  
$$G_{|B}=g,\quad \|G\|_{H^1(\R^d)} \le C_d \|g\|_{H^1(B)},\quad  \|G\|_{L^2(\R^d)} \le C_d \|g\|_{L^2(B)}.$$
Thus it suffices to show the Gagliardo-Nirenberg type inequality
\begin{align} \label{eq:G-H-U}
\|G\|_{L^{2(1+2/d)}(\R^d)} \le C_d \|G\|_{H^1(\R^d)}^{\frac d {d+2}} \|G\|_{L^2(\R^d)}^{\frac 2 {d+2}},\quad \forall G\in H^1(\R^d),
\end{align}
which is indeed equivalent to Sobolev's inequality $\|G\|_{L^{2(1+2/d)}} \le C_d \|G\|_{H^1}$ (Sobolev's inequality seems weaker, but we can apply it for $G_\ell(x)=\ell^{d/2} G(\ell x)$ and then optimize over $\ell>0$ to obtain \eqref{eq:G-H-U}). See also \cite[Lemma 8]{LunNamPor-16} for another derivation of \eqref{eq:G-H-U} via the fractional Sobolev embedding theorem. 
\end{proof}

\begin{lemma}[Local exclusion principle] \label{lem:LE}For any open ball $B\subset \R^d$ and any trace class operator $0\le \gamma \le 1$ on $L^2(B)$
\bq \label{eq:Neumann-gap}
\Tr(-\Delta_B \gamma) \ge \frac{1}{C_d |B|^{2/d}}  \Big( \int_B \rho_\gamma  - 1\Big).
\eq
\end{lemma}
\begin{proof} This bound goes back to Dyson and Lenard \cite{DysLen-67}. It is a consequence of the Neumann spectral gap 
$$
 -\Delta_B \ge  \frac{1}{C_d |B|^{2/d}} (\1_B- P)  \quad \text{ on }L^2(B)
$$
where $P= |B|^{-1} |\1_B\rangle\langle \1_B|$, the rank-one projection on the unique ground state $|B|^{-\frac 1 2} \1_B$ of the Neumann Laplacian. Then we can take the trace against $\gamma$ and conclude by using  $\Tr (\1_B\gamma)=\int_{B} \rho_\gamma$ and $\Tr (P\gamma)\le \Tr P =1$ (since $0\le \gamma \le 1$).  
\end{proof}

Now we are ready to provide 

\begin{proof}[Proof of Lemma \ref{lem:LT-ball}] Denote $M=\int_B \rho_\gamma$. From Lemmas \ref{lem:HO} and \ref{lem:LU} we have
$$
\Tr(-\Delta_B \gamma)  \ge \int_B |\nabla \sqrt{\rho_\gamma}|^2 \ge \frac{1}{C_d M^{2/d}}  \int_B \rho_\gamma^{1+2/d} - \frac{C_d}{ |B|^{2/d}} M.   
$$
Combining with \eqref{eq:Neumann-gap} we find that for every $\eps>0$, 
$$
 (1+\eps) \Tr(-\Delta_B \gamma)  \ge   \frac{\eps}{C_d M^{2/d}}  \int_B \rho_\gamma^{1+2/d}  + \frac{1}{ |B|^{2/d}} \Big[ \frac{M-1}{C_d} - \eps C_d M\Big].
$$
When $M=2$, we can always choose $\eps= \eps_d >0$ such that 
$$
\frac{M-1}{C_d} - \eps C_d M \ge 0,
$$
and the desired estimate follows. 
\end{proof}

\bigskip
\noindent
{\bf Step 2:} 
In order to put  Lemma \ref{lem:LT-ball} in good use, we need

\begin{lemma}[Besicovitch covering lemma \cite{Besicovitch-45,Morse-47}] Let $E$ be a bounded subset of $\R^d$. Let $\mathcal{F}$ be a collection of balls in $\R^d$ such that every point $x\in E$ is the center of a ball from $\mathcal{F}$. Then there is a sub-collection $\mathcal{G} \subset \mathcal{F}$ such that 
$$
\1_E \le \sum_{B\in \mathcal G} \1_B \le b_d,
$$
namely $E$ is covered by $\bigcup_{B\in \mathcal G} B$ and every point in $E$ belongs to at most $b_d$ balls from $\mathcal{G}$. Here the constant $b_d>0$  depends only on the dimension $d\ge 1$ (in particular, $b_d$ is independent of $E$).
\end{lemma}

Now we conclude 

\begin{proof}[Proof of the Lieb--Thirring inequality \eqref{eq:LT}] By a density argument it suffices to assume that $\gamma= \sum_{i=1}^N \lambda_n |u_n\rangle \langle u_n|$ with $N<\infty$, with orthonormal functions $\{u_n\}_{n=1}^N \subset C_c^\infty(\R^d)$, and with $0\le \lambda_n \le 1$. 

If $\int_{\R^d}\rho_\gamma\le 2$, then from Lemma \ref{lem:HO} and Sobolev's inequality \eqref{eq:S} we find that 
$$
\Tr (-\Delta \gamma) \ge \int_{\R^d} |\nabla \sqrt{\rho_\gamma}|^2 \ge \frac{\int_{\R^d} \rho_\gamma^{1+\frac 2 d}}{ C_d (\int_{\R^d} \rho_\gamma)^{2/d}} \ge \frac{1}{C_d 2^{\frac 2 d}}\int_{\R^d} \rho_\gamma^{1+\frac 2 d}. 
$$
Hence, it remains to consider the case 
$\int_{\R^d} \rho_\gamma \ge 2.$ Since $\rho_\gamma$ is continuous and its support $E\subset \R^d$ is bounded, for every $x\in E$ we can find a ball $B_x \subset \R^d$ centered at $x$ such that 
\bq  \label{eq:choice-B}
\int_{B_x} \rho_\gamma = 2. 
\eq
Applying the Besicovitch covering lemma, from the collection of balls $\mathcal{F}=\{B_x\}_{x\in E}$, we can find a sub-collection $\mathcal{G} \subset \mathcal {F}$ such that 
\bq  \label{eq:E-subset-G}
\1_E \le \sum_{B\in \mathcal G} \1_B \le b_d.
\eq
The second bound in \eqref{eq:E-subset-G} implies that
\begin{align*} 
b_d \int_{\R^d} |\nabla u_n|^2 \ge  \sum_{B\in \mathcal{G}} \int_{B} |\nabla u_n|^2
\end{align*} 
for all $n$. Therefore, 
\begin{align*} 
b_d \Tr (-\Delta \gamma) &=  \sum_{n=1}^N \lambda_n b_d \int_{\R^d} |\nabla u_n|^2 \ge   \sum_{n=1}^N  \lambda_n \sum_{B\in \mathcal{G}}  \int_{B} |\nabla u_n|^2\\
&= \sum_{B\in \mathcal{G}}  \sum_{n=1}^N \lambda_n  \int_{B} |\nabla u_n|^2   = \sum_{B \in \mathcal G} \Tr (-\Delta_{B} \gamma). 
\end{align*} 
On the other hand, for every ball $B \subset \mathcal G$, we have $\int_{B} \rho_\gamma = 2$ and $0\le \gamma \le 1$ on $L^2(B)$ (since $0\le \gamma \le 1$ on $L^2(\R^d)$). Thus Lemma \ref{lem:LT-ball} gives 
\bq 
\Tr(-\Delta_B \gamma) \ge \frac 1 C_d \int_{B} \rho_\gamma ^{1+\frac{2}{d}}.
\eq
Finally, combining the local bound \eqref{eq:E-subset-G} with the covering property \eqref{eq:E-subset-G} we conclude that
\bq 
b_d \Tr (-\Delta \gamma) \ge \sum_{B \in \mathcal G} \Tr (-\Delta_{B} \gamma) \ge  \frac 1 {C_d}  \sum_{B \in \mathcal G} \int_{B} \rho_\gamma^{1+2/d} \ge    \frac 1 {C_d} \int_{\R^d} \rho_\gamma^{1+2/d}.
\eq
The proof is complete. 
\end{proof}

\end{document}